# Comprehensive Review of Neural Differential Equations for Time Series Analysis

### v1

@@ 3.4 Stability and Convergence
NODEs and NCDEs achieve stability through Lipschitz constraints on vector fields (Haber and Ruthotto, 2017; Chen et al., 2018; Kidger et al., 2020). For NODEs, spectral norm conditions on network weights keep trajectories bounded (Kidger et al., 2021a). NCDEs maintain stability by bounding $\mathrm{d}X/\mathrm{d}t$ (Morrill et al., 2021a).
@@
NSDEs follow dynamics $\mathrm{d}{\bm{z}}(t)=f\bigl{(}{\bm{z}},t;\theta_{f}\bigr{)}\,\mathrm{d}t+g\bigl{(}{\bm{z}},t;\theta_{g}\bigr{)}\,\mathrm{d}W(t)$, where stability depends on drift $f$ and diffusion $g$ terms (Tzen and Raginsky, 2019; Kidger et al., 2021b). Oh et al. (2024) shows the stochastic stability of the proposed Neural SDEs (Langevin-type SDEs, linear-noise SDEs, and geometric SDEs) under suitable regularity conditions and relates the findings to their robustness against distribution shifts.
@@
The convergence of continuous-time methods can be analyzed through optimal transport theory, with the Wasserstein metric offering a principled measure of stability and generalization (Villani and others, 2009; Peyré et al., 2019). Formally, $(\Omega,\mathcal{F},P)$ specifies a probability space for random variables in $L^{p}(\Omega)$ whenever stochastic components arise. For an $\mathbb{R}^{d}$-valued random variable $X$, its law $\mathcal{L}(X)$ belongs to $\mathcal{P}(\mathbb{R}^{d})$. The Wasserstein distance of order $p$ measures how far two distributions $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$ are in terms of the minimum cost coupling, such as:
@@
Here, $\mathcal{C}(\mu,\nu)$ denotes the set of couplings whose marginals match $\mu$ and $\nu$. In the context of NDE models, particularly in stochastic or high-dimensional regimes, this metric provides a valuable tool for analyzing model robustness, generalization, and convergence properties (Kidger, 2021; Ruiz-Balet and Zuazua, 2023; Oh et al., 2024).
@@
NDEs rely on numerical integrators such as Euler, Runge–Kutta, and Dormand–Prince for deterministic models (Chen et al., 2018; Rubanova et al., 2019) and Euler–Maruyama, Milstein, or Reversible Heun for stochastic models (Kidger et al., 2021c; Tzen and Raginsky, 2019). On the other hand, systems that are stiff, exhibit jumps, or involve high-dimensional dynamics often necessitate specialized implicit solvers to ensure stability and computational efficiency (Jia and Benson, 2019; Rackauckas et al., 2020; Michoski et al., 2020; Kim et al., 2021).
@@ 4.1 Optimization of NDE-based Models
making it suitable for training models with terminal loss functions. However, the adjoint sensitivity method encounters numerical challenges, particularly in stiff or chaotic systems where reverse-time integration can amplify floating-point errors. Alternatively, checkpointing techniques (Gholami et al., 2019; Zhuang et al., 2020), store selected intermediate states during the forward pass, allowing localized recomputation during backward propagation.
@@
Beyond terminal loss functions, NDEs have been extended to incorporate integral loss functions distributed across the entire depth domain $S$. In the context of optimal control (Pontryagin, 2018), the integral loss is defined as formulated by Massaroli et al. (2020a):
@@
where the loss combines terminal contributions $\mathcal{L}(\bm{z}(S))$ with intermediate terms $\ell(t,\bm{z}(t))$. This formulation allows the latent state $\bm{z}(t)$ to evolve through a continuum of layers, guiding the model output toward the desired trajectory over the entire depth domain $S$. The adjoint dynamics for such integral loss functions are modified to include an additional term, $-\partial\ell/\partial\bm{z}(t)$, accounting for the distributed loss in the backward gradients (Grathwohl et al., 2018; Massaroli et al., 2020a; Finlay et al., 2020b). This approach enhances the control and flexibility of trajectory-level learning, enabling improved performance in tasks that require supervision over the entire temporal or spatial domain.
@@
For discrete methods like RNNs, it is a common optimization strategy, where the model’s recurrence is unfolded into a sequence of differentiable operations. However, these discrete methods face vanishing or exploding gradients over long sequences (Hochreiter and Schmidhuber, 1997; Pascanu et al., 2013; Goodfellow et al., 2016). In case of NDE-based variants, such as GRU-ODE (De Brouwer et al., 2019) and ODE-LSTM (Lechner and Hasani, 2020), mitigate these issues by replacing discrete recurrence with continuous evolution, enabling robust modeling of time series while retaining mechanisms like gating for temporal dependencies.
@@ 4.2 Regularization Methods
Standard regularization techniques have further refined NDE-based methods; stochastic sampling of end times (Ghosh et al., 2020), continuous-time dropout (Liu et al., 2020), and temporal adaptive batch normalization (Zheng et al., 2024). These methods form a comprehensive toolkit for advancing the performance and scalability of NDE-based models.
@@ 4.3 Numerical Solution of NDEs
Discrete-time methods, including variants of RNNs and Transformers, rely on layer-wise or self-attention updates without numeric integration. Although this design simplifies backpropagation and typically scales well, these models often struggle to accommodate irregular sampling unless augmented by masking or gating mechanisms (Chung et al., 2014). RNN-based architectures can also experience vanishing or exploding gradients for lengthy sequences unless they incorporate techniques such as gradient clipping or gated units (Hochreiter and Schmidhuber, 1997; Pascanu et al., 2013). Transformers replace recurrence with attention mechanisms but depend heavily on positional encodings and expansive parameter sets, which can limit their practicality for extensive time spans with variable sampling (Vaswani et al., 2017; Rae et al., 2019; Wen et al., 2022). By contrast, continuous-time NDEs use solvers that naturally adapt to asynchronous events, though solver overhead and potential stiffness require careful solver selection and parameter tuning.
@@
NDE-based models rely on numerical integration to evolve their latent states ${\bm{z}}(t)$ from initial conditions to final outputs. A fixed-step solver, such as explicit Euler or a basic Runge–Kutta method, updates ${\bm{z}}(t)$ at uniform intervals and is computationally simpler but may face instability if the dynamics change abruptly or if the time step is too large (Chen et al., 2018). Adaptive-step solvers, such as Dormand–Prince method, refine the step size $\Delta t$ in response to local error estimates, offering higher accuracy for stiff or highly non-uniform dynamics (Kidger et al., 2021c); however, this adaptive nature can cause unpredictable run times in large-scale tasks (Rackauckas et al., 2020). In the stochastic setting of NSDEs, solvers like Euler–Maruyama or Milstein handle the Brownian increments for the diffusion term, balancing efficiency and stability under well-defined noise (Tzen and Raginsky, 2019; Oh et al., 2024).
@@ 5 Comparison of NDE-based Methods
(Lechner and Hasani, 2020)
$&\bar{\bm{h}}_{t}=\text{ODESolve}\bigl{(}f,\bm{h}_{t-1},[t_{t-1},t_{t}]\bigr{)}\\ &\bm{h}_{t},\bm{c}_{t}=\text{LSTMCell}\bigl{(}\bm{c}_{t-1},\bar{\bm{h}}_{t},\bm{x}_{t}\bigr{)}\quad\text{where internal memory}\;\bm{c}_{t}\\ \end{aligned}$ Continuous-time modeling, Time series classification Synthetic data, Activity recognition, MNIST, MuJoCo physics
@@
(Kidger et al., 2020)
$&\mathrm{d}\bm{z}(t)=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t)\\ &\text{using Riemann–Stieltjes integral over control path}\;X(t)\end{aligned}$ Irregular time series analysis, Time series classification CharacterTrajectories, Speech Commands, PhysioNet Sepsis
@@
(Morrill et al., 2021a)
$&\mathrm{d}\bm{z}(t)=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)\\ &\text{where}\;\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)\;\text{represents the log-signature of}\;X(t)\end{aligned}$ Long time series analysis,   Robust feature extraction EigenWorms, BIDMC
@@
(Oh et al., 2025)
$&\frac{\mathrm{d}\bm{z}(t)}{\mathrm{d}t}=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t):=f^{*}(s;\theta_{f},X)\,{\mathrm{d}}t\\ &\hat{{\bm{z}}}(t)={\mathcal{G}}\bigl{(}t,\hat{{\bm{z}}}(0);\theta_{\mathcal{G}}\bigr{)}\quad\text{where}\;\frac{\mathrm{d}{\mathcal{G}}\bigl{(}t,\hat{{\bm{z}}}(0)\bigr{)}}{\mathrm{d}t}=f^{*}(t;\theta_{f},X)\end{aligned}$ Irregular time series classification, imputation, and forecasting UEA/UCR archive, PhysioNet mortality, PhysioNet Sepsis, Google stock, MuJoCo physics
@@
(Li et al., 2020)
$&\mathrm{d}\widetilde{\bm{z}}(t)=h_{\theta}\bigl{(}t,\widetilde{\bm{z}}(t)\bigr{)}\,\mathrm{d}t+\sigma\bigl{(}t,\widetilde{\bm{z}}(t)\bigr{)}\,\mathrm{d}W(t)\quad\;\text{(prior)}\\ &\mathrm{d}\bm{z}(t)=h_{\phi}\bigl{(}t,\bm{z}(t)\bigr{)}\,\mathrm{d}t+\sigma\bigl{(}t,\bm{z}(t)\bigr{)}\,\mathrm{d}W(t)\quad\;\text{(approximate posterior)}\\ \end{aligned}$ Stochastic adjoint sensitivity method Synthetic data, Motion capture
@@
(Oh et al., 2024)
$&\textbf{Neural LSDE:}\;{\mathrm{d}}{\bm{z}}(t)=\gamma\bigl{(}\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma})\,{\mathrm{d}}W(t)\\[1.39998pt] &\textbf{Neural LNSDE:}\;{\mathrm{d}}{\bm{z}}(t)=\gamma\bigl{(}t,\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\bm{z}}(t)\,{\mathrm{d}}W(t)\\[1.39998pt] &\textbf{Neural GSDE:}\;\frac{{\mathrm{d}}{\bm{z}}(t)}{{\bm{z}}(t)}=\gamma\bigl{(}t,\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma})\,{\mathrm{d}}W(t)\end{aligned}$ Irregular time series classification, imputation, and forecasting UEA/UCR archive, Speech Commands, PhysioNet mortality, PhysioNet Sepsis, MuJoCo physics
@@ 5.1 NODE Methods
Standard form of Neural ODE (Chen et al., 2018) introduced a novel approach to continuous-time modeling by parameterizing the hidden state dynamics with a neural network, allowing for the seamless handling of irregularly sampled time series data. Augmented ODE (Dupont et al., 2019) extend this framework by augmenting the state space with auxiliary variables, enabling the modeling of complex trajectories and avoiding trajectory overlap. ODE2VAE (Yildiz et al., 2019) combines NODEs with variational autoencoders to model latent dynamics in irregularly time series by learning both the continuous-time evolution and the underlying latent space.
On the other hand, GRU-ODE (De Brouwer et al., 2019) adopts continuous gating mechanisms inspired by GRU cells, where the hidden state evolves dynamically based on time gaps. $\bm{u}(\cdot)$, and $\bm{g}(\cdot)$ are continuous counterpart of GRU. ODE-RNN (Rubanova et al., 2019) combines ODE-based continuous evolution with discrete updates at observation points using an RNN cell. Similar to that, ODE-LSTM (Lechner and Hasani, 2020) incorporates the memory mechanisms of LSTM into the ODE framework, evolving the hidden state between observations using ODE solvers and performing discrete LSTM updates at observation points.

### v2

@@ 3.4 Stability and Convergence
NODEs and NCDEs achieve stability through Lipschitz constraints on vector fields Haber and Ruthotto (2017); Chen et al. (2018); Kidger et al. (2020). For NODEs, spectral norm conditions on network weights keep trajectories bounded Kidger et al. (2021a). NCDEs maintain stability by bounding $\mathrm{d}X/\mathrm{d}t$ Morrill et al. (2021a).
@@
NSDEs follow dynamics $\mathrm{d}{\bm{z}}(t)=f\bigl{(}{\bm{z}},t;\theta_{f}\bigr{)}\,\mathrm{d}t+g\bigl{(}{\bm{z}},t;\theta_{g}\bigr{)}\,\mathrm{d}W(t)$, where stability depends on drift $f$ and diffusion $g$ terms Tzen and Raginsky (2019); Kidger et al. (2021b). Oh et al. (2024) shows the stochastic stability of the proposed Neural SDEs (Langevin-type SDEs, linear-noise SDEs, and geometric SDEs) under suitable regularity conditions and relates the findings to their robustness against distribution shifts.
@@
The convergence of continuous-time methods can be analyzed through optimal transport theory, with the Wasserstein metric offering a principled measure of stability and generalization Villani and others (2009); Peyré et al. (2019). Formally, $(\Omega,\mathcal{F},P)$ specifies a probability space for random variables in $L^{p}(\Omega)$ whenever stochastic components arise. For an $\mathbb{R}^{d}$-valued random variable $X$, its law $\mathcal{L}(X)$ belongs to $\mathcal{P}(\mathbb{R}^{d})$. The Wasserstein distance of order $p$ measures how far two distributions $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$ are in terms of the minimum cost coupling, such as:
@@
Here, $\mathcal{C}(\mu,\nu)$ denotes the set of couplings whose marginals match $\mu$ and $\nu$. In the context of NDE models, particularly in stochastic or high-dimensional regimes, this metric provides a valuable tool for analyzing model robustness, generalization, and convergence properties Kidger (2021); Ruiz-Balet and Zuazua (2023); Oh et al. (2024).
@@
NDEs rely on numerical integrators such as Euler, Runge–Kutta, and Dormand–Prince for deterministic models Chen et al. (2018); Rubanova et al. (2019) and Euler–Maruyama, Milstein, or Reversible Heun for stochastic models Kidger et al. (2021c); Tzen and Raginsky (2019). On the other hand, systems that are stiff, exhibit jumps, or involve high-dimensional dynamics often necessitate specialized implicit solvers to ensure stability and computational efficiency Jia and Benson (2019); Rackauckas et al. (2020); Kim et al. (2021).
@@ 4.1 Optimization of NDE-based Models
making it suitable for training models with terminal loss functions. However, the adjoint sensitivity method encounters numerical challenges, particularly in stiff or chaotic systems where reverse-time integration can amplify floating-point errors. Alternatively, checkpointing techniques Gholami et al. (2019); Zhuang et al. (2020), store selected intermediate states during the forward pass, allowing localized recomputation during backward propagation.
@@
Beyond terminal loss functions, NDEs have been extended to incorporate integral loss functions distributed across the entire depth domain $S$. In the context of optimal control Pontryagin (2018), the integral loss is defined as formulated by Massaroli et al. (2020a):
@@
where the loss combines terminal contributions $\mathcal{L}(\bm{z}(S))$ with intermediate terms $\ell(t,\bm{z}(t))$. This formulation allows the latent state $\bm{z}(t)$ to evolve through a continuum of layers, guiding the model output toward the desired trajectory over the entire depth domain $S$. The adjoint dynamics for such integral loss functions are modified to include an additional term, $-\partial\ell/\partial\bm{z}(t)$, accounting for the distributed loss in the backward gradients Grathwohl et al. (2018); Massaroli et al. (2020a); Finlay et al. (2020b). This approach enhances the control and flexibility of trajectory-level learning, enabling improved performance in tasks that require supervision over the entire temporal or spatial domain.
@@
For discrete methods like RNNs, backpropagation is a common optimization strategy, where the model’s recurrence is unfolded into a sequence of differentiable operations. However, these discrete methods face vanishing or exploding gradients over long sequences Hochreiter and Schmidhuber (1997); Pascanu et al. (2013); Goodfellow et al. (2016). NDE-based variants, such as GRU-ODE De Brouwer et al. (2019), ODE-RNN Rubanova et al. (2019), ODE-LSTM Lechner and Hasani (2020), mitigate these issues by replacing discrete recurrence with continuous evolution, enabling robust modeling of time series while retaining mechanisms like gating for temporal dependencies.
@@ 4.2 Regularization Methods
Standard regularization techniques have further refined NDE-based methods; stochastic sampling of end times Ghosh et al. (2020), continuous-time dropout Liu et al. (2020), and temporal adaptive batch normalization Zheng et al. (2024). These methods form a comprehensive toolkit for advancing the performance and scalability of NDE-based models.
@@ 4.3 Numerical Solution of NDEs
Discrete-time methods, including variants of RNNs and Transformers, rely on layer-wise or self-attention updates without numeric integration. Although this design simplifies backpropagation and typically scales well, these models often struggle to accommodate irregular sampling unless augmented by masking or gating mechanisms Chung et al. (2014). RNN-based architectures can also experience vanishing or exploding gradients for lengthy sequences unless they incorporate techniques such as gradient clipping or gated units Hochreiter and Schmidhuber (1997); Pascanu et al. (2013). Transformers replace recurrence with attention mechanisms but depend heavily on positional encodings and expansive parameter sets, which can limit their practicality for extensive time spans with variable sampling Vaswani et al. (2017); Rae et al. (2019); Wen et al. (2023). By contrast, continuous-time NDEs use solvers that naturally adapt to asynchronous events, though solver overhead and potential stiffness require careful solver selection and parameter tuning.
@@
NDE-based models rely on numerical integration to evolve their latent states ${\bm{z}}(t)$ from initial conditions to final outputs. A fixed-step solver, such as explicit Euler or a basic Runge–Kutta method, updates ${\bm{z}}(t)$ at uniform intervals and is computationally simpler but may face instability if the dynamics change abruptly or if the time step is too large Chen et al. (2018). Adaptive-step solvers, such as Dormand–Prince method, refine the step size $\Delta t$ in response to local error estimates, offering higher accuracy for stiff or highly non-uniform dynamics Kidger et al. (2021c); however, this adaptive nature can cause unpredictable run times in large-scale tasks Rackauckas et al. (2020). In the stochastic setting of NSDEs, solvers like Euler–Maruyama or Milstein handle the Brownian increments for the diffusion term, balancing efficiency and stability under well-defined noise Tzen and Raginsky (2019); Oh et al. (2024).
@@ 5 Comparison of NDE-based Methods
ODE2VAE
Yildiz et al. (2019)
$&{\bm{z}}(t)=(\bm{s}(t),\bm{v}(t)),\quad\dot{\bm{s}}(t)=\bm{v}(t),\quad\dot{\bm{v}}(t)=f\bigl{(}\bm{s}(t),\bm{v}(t)\bigr{)},\\ &q_{\text{enc}}(\bm{z}(0)|\bm{x}_{0:N})=q_{\text{enc}}\left(\begin{pmatrix}\bm{s}(0)\\ \bm{v}(0)\end{pmatrix}\;\middle|\;\bm{x}_{0:N}\right)\;\text{(variational approximation)}\end{aligned}$ High-dimensional sequential data modeling, Motion prediction, Image sequence forecasting CMU Motion Capture, MNIST, Bouncing Balls
@@
Lechner and Hasani (2020)
$&\bar{\bm{h}}_{t}=\text{ODESolve}\bigl{(}f,\bm{h}_{t-1},[t_{t-1},t_{t}]\bigr{)}\\ &\bm{h}_{t},\bm{c}_{t}=\text{LSTMCell}\bigl{(}\bm{c}_{t-1},\bar{\bm{h}}_{t},\bm{x}_{t}\bigr{)},\quad\text{where internal memory}\;\bm{c}_{t}\\ \end{aligned}$ Continuous-time modeling, Time series classification Synthetic data, Activity recognition, MNIST, MuJoCo physics
@@
Kidger et al. (2020)
$&\mathrm{d}\bm{z}(t)=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t),\\ &\text{using Riemann–Stieltjes integral over control path}\;X(t)\end{aligned}$ Irregular time series analysis, Time series classification CharacterTrajectories, Speech Commands, PhysioNet Sepsis
@@
Morrill et al. (2021a)
$&\mathrm{d}\bm{z}(t)=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X),\\ &\text{where}\;\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)\;\text{represents the log-signature of}\;X(t)\end{aligned}$ Long time series analysis,   Robust feature extraction EigenWorms, BIDMC
Neural Lad
Li et al. (2023)
[…]
@@
Oh et al. (2025)
$&\frac{\mathrm{d}\bm{z}(t)}{\mathrm{d}t}=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t):=f^{*}(t;\theta_{f},X)\,{\mathrm{d}}t\\ &\hat{{\bm{z}}}(t)={\mathcal{G}}\bigl{(}t,\hat{{\bm{z}}}(0);\theta_{\mathcal{G}}\bigr{)}\quad\text{where}\;\frac{\mathrm{d}{\mathcal{G}}\bigl{(}t,\hat{{\bm{z}}}(0)\bigr{)}}{\mathrm{d}t}=f^{*}(t;\theta_{f},X)\end{aligned}$ Irregular time series classification, imputation, and forecasting UEA/UCR archive, PhysioNet mortality, PhysioNet Sepsis, Google stock, MuJoCo physics
@@
Li et al. (2020)
$&\mathrm{d}\widetilde{\bm{z}}(t)=h_{\theta}\bigl{(}t,\widetilde{\bm{z}}(t)\bigr{)}\,\mathrm{d}t+\sigma\bigl{(}t,\widetilde{\bm{z}}(t)\bigr{)}\,\mathrm{d}W(t)\quad\;\text{(prior)}\\ &\mathrm{d}\bm{z}(t)=h_{\phi}\bigl{(}t,\bm{z}(t)\bigr{)}\,\mathrm{d}t+\sigma\bigl{(}t,\bm{z}(t)\bigr{)}\,\mathrm{d}W(t)\quad\;\text{(approximate posterior)}\\ \end{aligned}$ Stochastic adjoint sensitivity method Synthetic data, Motion capture data
@@
Oh et al. (2024)
$&\textbf{Neural LSDE:}&{\mathrm{d}}{\bm{z}}(t)&=\gamma\bigl{(}\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma})\,{\mathrm{d}}W(t)\\[1.39998pt] &\textbf{Neural LNSDE:}&{\mathrm{d}}{\bm{z}}(t)&=\gamma\bigl{(}t,\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\bm{z}}(t)\,{\mathrm{d}}W(t)\\[1.39998pt] &\textbf{Neural GSDE:}&\frac{{\mathrm{d}}{\bm{z}}(t)}{{\bm{z}}(t)}&=\gamma\bigl{(}t,\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma})\,{\mathrm{d}}W(t)\end{aligned}$ Irregular time series classification, imputation, and forecasting UEA/UCR archive, Speech Commands, PhysioNet mortality, PhysioNet Sepsis, MuJoCo physics
@@ 5.1 NODE Methods
Standard form of Neural ODE Chen et al. (2018) introduced a novel approach to continuous-time modeling by parameterizing the hidden state dynamics with a neural network, allowing for the seamless handling of irregularly sampled time series data. Augmented ODE Dupont et al. (2019) extend this framework by augmenting the state space with auxiliary variables, enabling the modeling of complex trajectories and avoiding trajectory overlap. ODE2VAE Yildiz et al. (2019) combines NODEs using Bayesian Neural Network (BNN) with variational autoencoders to model latent dynamics in irregular time series by learning both the continuous-time evolution and the underlying latent space.
On the other hand, GRU-ODE De Brouwer et al. (2019) adopts continuous gating mechanisms inspired by GRU cells, where the hidden state evolves dynamically based on time gaps. $\bm{u}(\cdot)$, and $\bm{g}(\cdot)$ are continuous counterparts of GRU. ODE-RNN Rubanova et al. (2019) combines ODE-based continuous evolution with discrete updates at observation points using an RNN cell. Similar to that, ODE-LSTM Lechner and Hasani (2020) incorporates the memory mechanisms of LSTM into the ODE framework, evolving the hidden state between observations using ODE solvers and performing discrete LSTM updates at observation points.